OpenStudy (anonymous):

find the slope of the line that goes through (-2,-6), "(-9,-17)

OpenStudy (anonymous):

i got this! do you know the slope formula?

OpenStudy (anonymous):

answer is 11 over 7

OpenStudy (anonymous):

yes, kinda hard to type

OpenStudy (anonymous):

thanks for your help alex

OpenStudy (anonymous):

yea, its y2-y1 over x2-x1, and no problem!

OpenStudy (bradely):

m = y2-y1/x2-x1 =(-17-(-6))/(-9-(-2)) =-11/-7 =11/7 Source: http://www.mathskey.com/question2answer/